OpenStudy (anonymous):

How much of an alloy with 30% copper must be added to 25 pounds of a second alloy that is 20% copper to create a 27% alloy?

jimthompson5910 (jim_thompson5910):

Let x = amount of the 30% copper alloy (in pounds) So if you have x pounds of the 30% copper alloy, then you will have 0.30x pounds of pure copper. If you have 25 pounds of a 20% copper alloy, then you will have 25*0.20 = 5 pounds of pure copper Combine the two to get a total of 0.30x + 5 pounds of pure copper

jimthompson5910 (jim_thompson5910):

There are a total of x+25 pounds of this alloy. We want the final amount to be a 27% copper alloy, so (amount of pure copper)/(amount of metal total) = desired percentage (0.30x + 5)/(x+25) = 0.27 solve that for x to get your answer